# Planning with Adaptive World Models for Autonomous Driving

### v1

@@ Abstract
Motion planning is crucial for safe navigation in complex urban environments. Historically, motion planners (MPs) have been evaluated with procedurally-generated simulators like CARLA. However, such synthetic benchmarks do not capture real-world multi-agent interactions. nuPlan, a recently released MP benchmark, addresses this limitation by augmenting real-world driving logs with closed-loop simulation logic, effectively turning the fixed dataset into a reactive simulator. We analyze the characteristics of nuPlan’s recorded logs and find that each city has its own unique driving behaviors, suggesting that robust planners must adapt to different environments. We learn to model such unique behaviors with BehaviorNet, a graph convolutional neural network (GCNN) that predicts reactive agent behaviors using features derived from recently-observed agent histories; intuitively, some aggressive agents may tailgate lead vehicles, while others may not. To model such phenomena, BehaviorNet predicts parameters of an agent’s motion controller rather than predicting its spacetime trajectory (as most forecasters do). Finally, we present AdaptiveDriver, a model-predictive control (MPC) based planner that unrolls different world models conditioned on BehaviorNet’s predictions. Our extensive experiments demonstrate that AdaptiveDriver achieves state-of-the-art results on the nuPlan closed-loop planning benchmark, reducing test error from 6.4% to 4.6%, even when applied to never-before-seen cities. We include additional videos on our project page.
Keywords: Motion Planning, Autonomous Vehicles, Data-Driven Simulation
1 Introduction
@@
Planning Evaluation. Although nuPlan evaluates algorithms on ego-centric forecasting accuracy (defined as the C1 metric  [3]) and motion planning in a non-reactive world (C2), our work primarily focuses on motion planning in a reactive world (C3) as this most closely resembles real-world deployment. Recent methods like PDM-C [4] have dramatically improved planning performance in reactive environments by combining model-predictive control (MPC) with classical motion controllers like the Intelligent Driver Model (IDM) [5].
@@
Different World Models for Each City. Intuitively, PDM-C could achieve perfect planning performance given an accurate world model of future agent actions and an infinite number of trajectory proposals. However, evaluating complex world models can be computationally challenging. Instead, PDM-C makes use of a simpler “world-on-rails” model where agents move with constant velocity during rollouts. Notably, this non-reactive model does not accurately simulate agent behaviors, leading to poorly scored proposals. To improve this “world-on-rails” model, we analyze the characteristics of nuPlan’s recorded logs and find that each city has its own unique driving behaviors. For example, Boston drivers are more aggressive (e.g. drive with lower average min-gap) than Pittsburgh drivers (cf. Fig. 1). We model such unique behaviors with BehaviorNet, a graph convolution-based network that predicts future agent actions based on features derived from recently observed agent behaviors. Unlike traditional forecasters, BehaviorNet parameterizes future agent behaviors with IDM controls to create a reactive world model. Conceptually, each set of predicted IDM control parameters unrolls a different world model.
Building an Adaptive Planner. Finally, we introduce AdaptiveDriver, an MPC-based planner that scores spatio-temporal proposals conditioned on BehaviorNet’s reactive world model. Extensive experiments demonstrate that our AdaptiveDriver planner achieves state-of-the-art results on the nuPlan closed-loop planning benchmark, reducing relative test error by 37.5% (from 6.4% to 4.6% absolute error). Importantly, the novelty of our work is in our parameterization of future agent behaviors as IDM control parameters (rather than non-reactive spacetime forecasts), which implicitly encodes an ensemble of reactive world models that generalize to never-before-seen environments.
@@ 2 Related Works
Rule-Based Planning. Although recent works focus on learning robust policies by predicting goal-conditioned way-points [6], cost-volumes, and reward functions [7], rule-based planners still outperform learning-based approaches on real data [4]. Rule-based planners are well studied  [8, 9, 10, 11, 12], and have been widely adopted due to their safety guarantees and interpretability [13, 14, 15, 16, 17]. Given the current position, velocity, and distance to the lead vehicle, rule-based planners estimate longitudinal acceleration to safely progress towards the target. The Intelligent Driver Model (IDM) [5] is a classic non-learned algorithm for vehicle motion planning that relies on graph-based search to reach the target while employing a PID velocity controller to avoid collisions with other vehicles. Dauner et. al. [4] upgrades IDM by sampling multiple trajectories and unrolling a constant velocity world model to select an optimal trajectory with minimum cost. This improves collision avoidance without resulting in overly conservative motion plans.
@@
Trajectory Optimization. Motion planning is often framed as an optimization problem of hand-designed cost functions, which are then minimized to generate an optimal trajectory [18, 19, 20, 21]. To simplify this process, cost functions assume a quadratic objective function or divide the planning task into its lateral and longitudinal components. Approaches such as A* [22], RRT [23], and dynamic programming [20] are commonly used to search for optimal solutions. CoverNet [24] generates a set of trajectories and evaluates them based on cost functions, selecting the trajectory with lowest cost. While these methods are attractive due to their parallelizability, interpretability and functional guarantees, they are not robust when applied to real-world scenarios and require significant hyperparameter tuning. Conventional trajectory optimization approaches typically aim to compute a complete trajectory that spans from the initial configuration to the desired goal configuration. However, given the inherently dynamic and uncertain nature of the driving environment, precise long-horizon motion plans cannot be predicted in advance. As a result, model-predictive control (MPC) has gained prominence in recent years for real-time path planning [25, 26, 27, 28] because it adopts an iterative cost minimization strategy to select a locally optimal trajectory for each timestep. This allows MPC-based algorithms to quickly adapt to changes in the environment.
@@
Data-Driven Simulation. In recent years, many learning-based planners have emerged, leveraging the availability of simulator environments like CARLA [1], AirSim [2], and others. However, current simulators are limited because they rely on synthetic data generated from game engines and have insufficient visual fidelity. Importantly, they lack the necessary diversity of driving scenarios required for comprehensive training and evaluation. To address these limitations, [29] proposes multi-agent behavior models for generating diverse and realistic traffic simulations. More recently, [30] introduced the Waymo sim agents challenge which evaluates simulators by comparing the trajectories of all simulated agents against their ground-truth trajectories. Furthermore, CommonRoad [31] offers a driving dataset and planning benchmark which combines real-world data and rule-based heuristics. In contrast, nuPlan [3] augments real-world driving logs with closed-loop simulation logic, effectively turning the fixed dataset into a reactive simulator. nuPlan has released $1300$ hours of real world driving logs from various cities including Las Vegas, Boston, Pittsburgh, and Singapore. Driving in each city presents a unique set of driving challenges. For example, Las Vegas has many high density pick-up and drop-off locations, and intersections with 8 parallel driving lanes per direction. In Boston, drivers tend to double park, creating distinct planning challenges.
3 Planning With An Ensemble of World Models
@@
Table 1: Optimized City-Specific Reactive World Models. We generate city-specific IDM parameters by optimizing the objective function described in Equation 1 over all training logs from a given city. The optimized world model parameters attempt to minimize the distribution shift between simulated trajectory rollouts and recorded logs while maximizing nuPlan’s C3 metric on the training set.
@@
Understanding PDM-C’s Limitations. PDM-C is a state-of-the-art rule-based planner that improves upon the Intelligent Driver Model (IDM), a car following algorithm that employs a simple longitudinal PID velocity controller along a reference path. PDM-C upgrades IDM to an MPC-based planner by modulating IDM’s reference path with different longitudinal velocities and lateral offsets to generate candidate trajectories, internally unrolling a world model for other agents, and selecting the trajectory that minimizes a cost function over that world model. Notably, PDM-C makes use of a simpler “world-on-rails” internal world model where other agents are non-reactive and move with constant velocity during rollouts. Although a ‘world-on-rails” model with constant velocity forecasts may work well for short-horizon forecasting, it fails to correctly simulate multi-agent interactions like lane changes, lane merges and stopping at traffic lights.
Predicting Future Agent Actions with BehaviorNet. We improve upon the “world-on-rails” model by learning to predict future agent behaviors with BehaviorNet. We model the unique driving characteristics of each scenario by encoding a vectorized road graph of radius $R$ around the ego-vehicle and two seconds of trajectory history for all nearby agents. BehaviorNet consists of several multi-scale graph convolution and attention modules followed by a fully connected layer to predict IDM control parameters. We further describe BehaviorNet’s architecture in the supplement. Notably, unlike traditional forecasters, BehaviorNet directly predicts IDM control parameters, which can then used to unroll a reactive world model.
@@
where $X_{SIM}$ is the position of all agents in the simulated IDM rollout and $X_{LOG}$ is the position of all agents in the recorded driving log. We assume that the ego-vehicle is non-reactive and simply replays its driving log. Intuitively, our optimization finds controls parameters that best tracks ground-truth agent positions. Importantly, Eq 1 could be optimized over individual agents, over individual logs, or over all logs from a particular city. Motivated by our observation on city-specific behaviors, we first present results for optimizing target IDM parameters for all logs from a particular city (cf. Table 1). We highlight salient observations below. First, $\boldsymbol{\theta}^{BOS}$ and $\theta^{LAS}$ tailgate vehicles more aggressively than $\boldsymbol{\theta}^{PIT}$ and $\boldsymbol{\theta}^{SIN}$, which better tracks the city-specific min-gap statistics from Figure 1. Second, simulated agents using the default IDM world model ($\boldsymbol{\theta}^{IDM}$, optimized to minimize Eq. 1 over the entire dataset) drive less aggressively with lower maximum acceleration and maximum deceleration than city-specific models from PIT ($\boldsymbol{\theta}^{PIT}$) and BOS ($\boldsymbol{\theta}^{BOS}$). Third, $\boldsymbol{\theta}^{BOS}$ seems to have the most aggressive drivers and $\boldsymbol{\theta}^{LAS}$ seems to have least aggressive ones, as defined by maximum acceleration and maximum deceleration.
@@
Figure 3: Visualizing Clusters of Behaviors. In (a), we optimize scenario-specific world models using Equation 1 and visualize per-log IDM parameters using a tSNE plot, coloring cities differently. For comparison, we also plot per-city-optimized IDM parameters (Table 1) as large colored dots. In (b), we cluster per-scenario IDM parameters using $K$-means and visualize different clusters. Each cluster represents a unique emergent driving behavior. In (c), we compare the distribution of min-gap between the “aggressive driver” cluster (left) and “passive driver” cluster (right). We note that the “aggressive driver” cluster has a lower average min-gap than the “passive driver” cluster, validating that our $K$-means clusters encode unique city-agnostic driving behaviors.
@@
Training Log-BehaviorNet. Although each city has distinct driving characteristics, agents can still behave differently within a city. For example, Boston drivers may tailgate within the city but drive more cautiously on highways. To model this, we simply optimize Eq.1 over each individual training log. Figure 3 (a) visualizes the set of log-specific IDM parameters $\{\boldsymbol{\theta}\}$ with tSNE, color-coded by city. Rather than training BehaviorNet to directly regress these parameters, we reframe the problem as a simpler discrete classification task. Specifically, we cluster the set of $\{\boldsymbol{\theta}\}$ into $K$ clusters and train BehaviorNet with a $K$-way softmax loss. We refer to this network as Log-BehaviorNet, in contrast to City-BehaviorNet. Figure 3-(b) compares the learned behavior clusters with the original city “clusters” in (a). Figure 3-(c) plots the min-gap distribution from two distinct clusters, suggesting that each cluster loosely corresponds to prototypical behaviors such as “aggressive” or “passive”. We tune the number of behavior clusters ($K$) so as to maximize $C_{3}$ performance on the nuPlan val-set. Interestingly, the optimal number (16) is far more than the number of distinct cities (4). Importantly, we show that city-agnostic clusters generalize better than city-specific models, particularly when evaluated on never-before-seen cities.
@@
Incorporating Learned Priors into Rule-Based Planners. Although rule-based planners like PDM-C still outperform learning-based methods on real data, they fail to accurately model future agent behaviors in world model rollouts. We aim to bridge the gap between rule-based and learning-based planners with AdaptiveDriver (cf. Fig. 4), a model-predictive control (MPC) planner that uses behavior parameter predictions to improve the quality of world model rollouts. Notably, although AdaptiveDriver and PDM-C are both instantations of MPC-based planners, our model (1) uses a reactive world-model that (2) adapts to each log using features derived from past agent behaviors.
       Model        City        C2        C3        Raster Model  [32]        All        78.50        75.67        LaneGCN [33]        All        78.52        77.12        UrbanDriver [7]        All        80.79        81.16        IDM [5]        All        83.57        81.61        PDM-C [4]        All        94.49        93.61        AdaptiveDriver w/ City-BehaviorNet (Ours)        All        94.97        95.13        AdaptiveDriver w/ Log-BehaviorNet (Ours)        All        95.15        95.35
@@
Table 2: nuPlan Benchmark Results. We evaluate several rule-based and learning-based planners on the nuPlan test set. First, we note that rule-based planners like IDM and PDM-C outperform prior learning-based methods like Raster Model, LaneGCN, and UrbanDriver on C2 and C3. In contrast, AdaptiveDriver demonstrates that carefully combining learning-based behavior priors with rule-based planners significantly improves over prior work, reducing test error from $6.4\%$ to $4.6\%$. Further, we find that log-specific behavior models (Log-BehaviorNet) slightly outperform city-specific behavior models (City-BehaviorNet). We provide additional visuals comparing our planner to PDM-C in Fig 5.
4 Experiments

### v3

@@ Abstract
Motion planning is crucial for safe navigation in complex urban environments. Historically, motion planners (MPs) have been evaluated with procedurally-generated simulators like CARLA. However, such synthetic benchmarks do not capture real-world multi-agent interactions. nuPlan, a recently released MP benchmark, addresses this limitation by augmenting real-world driving logs with closed-loop simulation logic, effectively turning the fixed dataset into a reactive simulator. We analyze the characteristics of nuPlan’s recorded logs and find that each city has its own unique driving behaviors, suggesting that robust planners must adapt to different environments. We learn to model such unique behaviors with BehaviorNet, a graph convolutional neural network (GCNN) that predicts reactive agent behaviors using features derived from recently-observed agent histories; intuitively, some aggressive agents may tailgate lead vehicles, while others may not. To model such phenomena, BehaviorNet predicts the parameters of an agent’s motion controller rather than directly predicting its spacetime trajectory (as most forecasters do). Finally, we present AdaptiveDriver, a model-predictive control (MPC) based planner that unrolls different world models conditioned on BehaviorNet’s predictions. Our extensive experiments demonstrate that AdaptiveDriver achieves state-of-the-art results on the nuPlan closed-loop planning benchmark, improving over prior work by 2% on Test-14 Hard R-CLS, and generalizes even when evaluated on never-before-seen cities.
I Introduction
@@
Planning Evaluation. Although nuPlan evaluates algorithms on ego-centric forecasting accuracy (defined as the OLS metric  [3]) and motion planning in a non-reactive world (NR-CLS), our work primarily focuses on motion planning in a reactive world (R-CLS) as this most closely resembles real-world deployment. Recent methods like PDM-C [4] have dramatically improved planning performance in reactive environments by combining model-predictive control (MPC) with classical motion controllers like the Intelligent Driver Model (IDM) [5].
@@
Different World Models for Each City. Intuitively, PDM-C could achieve perfect planning performance given an accurate world model of future agent actions and an infinite number of trajectory proposals. However, evaluating complex world models can be computationally infeasible. Instead, PDM-C makes use of a simpler “world-on-rails” model where agents move with constant velocity during rollouts. Notably, this non-reactive model does not accurately simulate agent behaviors, leading to poorly scored proposals. To improve this “world-on-rails” model, we analyze the characteristics of nuPlan’s recorded logs and find that each city has its own unique driving behaviors. For example, Boston drivers are more aggressive (e.g. drive with lower average min-gap) than Pittsburgh drivers (cf. Fig. 1). We model such unique behaviors with BehaviorNet, a graph convolution-based network that predicts future agent actions based on features derived from recently observed agent behaviors. Unlike traditional forecasters, BehaviorNet parameterizes future agent behaviors with IDM controls to create a reactive world model. Conceptually, each set of predicted control parameters unrolls a different world model.
Building an Adaptive Planner. Finally, we introduce AdaptiveDriver, an MPC-based planner that scores spatio-temporal proposals conditioned on BehaviorNet’s reactive world model. Extensive experiments demonstrate that our AdaptiveDriver planner achieves state-of-the-art results on the nuPlan closed-loop planning benchmark, improving by 2% on Test-14 Hard. Importantly, the novelty of our work is in our parameterization of future agent behaviors as control parameters (rather than non-reactive spacetime forecasts), which implicitly encodes an ensemble of reactive world models that generalize to never-before-seen environments.
@@ II Related Works
Rule-Based Planning. Although recent works focus on learning robust policies by predicting goal-conditioned way-points [6], cost-volumes, and reward functions [7], rule-based planners still outperform learning-based approaches on real data [4]. Rule-based planners are well studied  [8, 9, 10, 11, 12], and have been widely adopted due to their safety guarantees and interpretability [13, 14, 15, 16, 17]. Given the current position, velocity, and distance to the lead vehicle, rule-based planners estimate longitudinal acceleration to safely progress towards the target. The Intelligent Driver Model (IDM) [5] is a classic non-learned algorithm for vehicle motion planning that relies on graph-based search to reach the target while employing a PID velocity controller to avoid collisions with other vehicles. Dauner et. al. [4] upgrades IDM by sampling multiple trajectories and unrolling a constant velocity world model to select an optimal trajectory with minimum cost.
@@
Trajectory Optimization. Motion planning is often framed as an optimization problem of hand-designed cost functions, which are then minimized to generate an optimal trajectory [18, 19, 20, 21]. To simplify this process, cost functions assume a quadratic objective function or divide the planning task into its lateral and longitudinal components. Approaches such as A* [22], RRT [23], and dynamic programming [20] are commonly used to search for optimal solutions. CoverNet [24] generates a set of trajectories and evaluates them based on cost functions, selecting the trajectory with lowest cost. While these methods are attractive due to their parallelizability, interpretability and functional guarantees, they are not robust when applied to real-world scenarios and require significant hyperparameter tuning. Conventional trajectory optimization approaches typically aim to compute a complete trajectory that spans from the initial configuration to the desired goal configuration. However, given the inherently dynamic and uncertain nature of the driving environment, precise long-horizon motion plans cannot be predicted in advance. As a result, model-predictive control (MPC) has gained prominence in recent years for real-time path planning [25, 26, 27, 28] because it adopts an iterative cost minimization strategy to select a locally optimal trajectory for each timestep.
@@
Data-Driven Simulation. In recent years, many learning-based planners have emerged, leveraging the availability of simulator environments like CARLA [1], AirSim [2], and others. However, current simulators are limited because they rely on synthetic data generated from game engines and have insufficient visual fidelity. Importantly, they lack the necessary diversity of driving scenarios required for comprehensive training and evaluation. To address these limitations, [29] proposes multi-agent behavior models for generating diverse and realistic traffic simulations. More recently, [30] introduced the Waymo sim agents challenge which evaluates simulators by comparing the trajectories of all simulated agents against their ground-truth trajectories. Furthermore, CommonRoad [31] offers a driving dataset and planning benchmark which combines real-world data and rule-based heuristics. In contrast, nuPlan [3] augments real-world driving logs with closed-loop simulation logic, effectively turning the fixed dataset into a reactive simulator. nuPlan has released $1300$ hours of real world driving logs from various cities including Las Vegas, Boston, Pittsburgh, and Singapore. Driving in each city presents a unique set of challenges. For example, Las Vegas has many high density pick-up and drop-off locations, and intersections with 8 parallel driving lanes per direction. In contrast, Boston drivers tend to double park, creating distinct planning challenges.
III Planning With An Ensemble of Models
@@
TABLE I: Optimized City-Specific Reactive World Models. We generate city-specific IDM parameters ($\theta_{0},\dots,\theta_{4}$) by optimizing the objective function described in Equation 1 over all training logs from a given city. The optimized world model parameters attempt to minimize the distribution shift between simulated trajectory rollouts and recorded logs on the training set.
@@
Understanding PDM-C’s Limitations. PDM-C is a state-of-the-art rule-based planner that improves upon the Intelligent Driver Model (IDM), a car following algorithm that employs a simple longitudinal PID velocity controller along a reference path. PDM-C upgrades IDM to an MPC-based planner by modulating IDM’s reference path with different longitudinal velocities and lateral offsets to generate candidate trajectories, internally unrolling a world model for other agents, and selecting the trajectory that minimizes a cost function over that world model. Notably, PDM-C makes use of a simpler “world-on-rails” internal world model where other agents are non-reactive and move with constant velocity during rollouts. Although a ‘world-on-rails” model with constant velocity forecasts may work well for short-horizon forecasting, it fails to correctly simulate multi-agent interactions like lane changes and lane merges.
Predicting Future Agent Actions with BehaviorNet. We improve upon PDM-C’s “world-on-rails” model by learning to predict future agent behaviors with BehaviorNet. We model the unique driving characteristics of each scenario by encoding a vectorized road graph of radius $R$ around the ego-vehicle and two seconds of trajectory history for all nearby agents. BehaviorNet consists of several multi-scale graph convolution and attention modules followed by a fully connected layer to predict IDM control parameters. We further describe BehaviorNet’s architecture in Figure 3. Notably, unlike traditional forecasters, BehaviorNet directly predicts IDM control parameters ($\theta_{0},\dots,\theta_{4}$), which can then used to unroll a reactive world model.
@@
where $X_{SIM}$ is the position of all agents in the simulated IDM rollout and $X_{LOG}$ is the position of all agents in the recorded driving log. We assume that the ego-vehicle is non-reactive and simply replays its driving log. Intuitively, our optimization finds controls parameters that best tracks ground-truth agent positions. Importantly, Eq 1 could be optimized over individual agents, over individual logs, or over all logs from a particular city. Motivated by our observation on city-specific behaviors, we first present results for optimizing target IDM parameters for all logs from a particular city (cf. Table I). We highlight salient observations below. First, $\boldsymbol{\theta}^{BOS}$ and $\theta^{LAS}$ tailgate vehicles more aggressively than $\boldsymbol{\theta}^{PIT}$ and $\boldsymbol{\theta}^{SIN}$, which better tracks the city-specific min-gap statistics from Figure 1. Second, simulated agents using the default IDM world model ($\boldsymbol{\theta}^{IDM}$, optimized to minimize Eq. 1 over the entire dataset) drive less aggressively with lower maximum acceleration and maximum deceleration than city-specific models from PIT ($\boldsymbol{\theta}^{PIT}$) and BOS ($\boldsymbol{\theta}^{BOS}$). Third, $\boldsymbol{\theta}^{BOS}$ seems to have the most aggressive drivers and $\boldsymbol{\theta}^{LAS}$ seems to have least aggressive ones, as defined by maximum acceleration and maximum deceleration.
@@
Figure 4: Visualizing Clusters of Behaviors. In (a), we optimize scenario-specific world models using Equation 1 and visualize per-log IDM parameters using a tSNE plot, coloring cities differently. For comparison, we also plot per-city-optimized IDM parameters (Table I) as large colored dots. In (b), we cluster per-scenario IDM parameters using $K$-means and visualize different clusters. Each cluster represents a unique emergent driving behavior. In (c), we compare the distribution of min-gap between the “aggressive driver” cluster (left) and “passive driver” cluster (right). We note that the “aggressive driver” cluster has a lower average min-gap than the “passive driver” cluster, validating that our $K$-means clusters encode unique city-agnostic driving behaviors.
@@
Training Log-BehaviorNet. Although each city has distinct driving characteristics, agents can still behave differently within a city. For example, Boston drivers may tailgate within the city but drive more cautiously on highways. To model this, we simply optimize Eq.1 over each individual training log. Figure 4 (a) visualizes the set of log-specific IDM parameters $\{\boldsymbol{\theta}\}$ with tSNE, color-coded by city. Rather than training BehaviorNet to directly regress these parameters, we reframe the problem as a simpler discrete classification task. Specifically, we cluster the set of $\{\boldsymbol{\theta}\}$ from the train-set into $K$ clusters and train BehaviorNet with a $K$-way softmax loss. We refer to this network as Log-BehaviorNet, in contrast to City-BehaviorNet. Figure 4-(b) compares the learned behavior clusters with the original city “clusters” in (a). Figure 4-(c) plots the min-gap distribution from two distinct clusters, suggesting that each cluster loosely corresponds to prototypical behaviors such as “aggressive” or “passive”. We tune the number of behavior clusters ($K$) so as to maximize $R-CLS$ performance on the nuPlan train-set. Interestingly, the optimal number (16) is far more than the number of distinct cities (4). Importantly, we show that city-agnostic clusters generalize better than city-specific models, particularly when evaluated on never-before-seen cities (cf. Table V).
@@
Adding Learned Priors into Rule-Based Planners. Although rule-based planners like PDM-C still outperform learning-based methods on real data, they fail to accurately model future agent behaviors in world model rollouts. We aim to bridge the gap between rule-based and learning-based planners with AdaptiveDriver (cf. Fig. 5), a model-predictive control (MPC) planner that uses behavior parameter predictions to improve the quality of world model rollouts. Notably, although AdaptiveDriver and PDM-C are both instantations of MPC-based planners, our model improves upon PDM-C using (1) a reactive world-model that (2) adapts to each log using features derived from past agent behaviors.
       Val-14        Test-14 Hard        Model        NR-CLS        R-CLS        NR-CLS        R-CLS        Log Replay        94.03        75.86        85.96        68.80        IDM  [5]        70.39        72.42        56.16        62.26        PDM-C  [4]        92.51        91.79        65.07        75.18        RasterModel  [33]        69.66        67.54        49.47        52.16        UrbanDriver  [7]        63.27        61.02        51.54        49.07        GC-PGP  [34]        55.99        51.39        43.22        39.63        PDM-O  [4]        52.80        57.23        33.51        35.83        GameFormer  [35]        80.80        79.31        66.59        68.83        PlanTF  [36]        84.83        76.78        72.68        61.70        DTPP  [37]        89.64        89.78        59.44        62.94        LLM-AssistUNC  [38]        90.11        90.32        -        -        LLM-AssistPAR  [38]        93.05        92.20        -        -        PlanAgent  [39]        93.26        92.75        72.51        76.82        AdaptiveDriver w/ City-BehaviorNet (Ours)        92.97        93.28        70.11        78.02        AdaptiveDriver w/ Log-BehaviorNet (Ours)        93.15        93.49        73.25        78.87
@@
TABLE II: nuPlan Benchmark Results. We evaluate several rule-based and learning-based planners on the nuPlan test set. First, we note that rule-based planners like IDM and PDM-C outperform prior learning-based methods like Raster Model, LaneGCN, and UrbanDriver on NR-CLS and R-CLS. In contrast, AdaptiveDriver demonstrates that carefully combining learning-based behavior priors with rule-based planners significantly improves over prior work by 2% on Test14-Hard R-CLS. Results of prior work are reproduced from [36]. We provide additional visuals comparing our planner to our PDM-C baseline in Fig 6.
IV Experiments